Controls Meeting 8-Jan-2015
- Work on the gas system PLC (Dave)
- Remote reset for CAENs (Dave, Werth)
- Warm magnet ramping procedures (Hovanes, Mark)
- Controls work on solenoid (Yi, Mark)
- Power supply clock settings (Mark)
- Missing controls for detector voltages (Hovanes)
- Adding more alarms ? (Hovanes, Yi)
- Any other business
- New motors
9:30AM Thurs 8-January-2015 CC A110
Present: Dave, Mark, Scot, Hovanes
- Yi will be leaving at the end of January. Hall D group needs to learn the slow controls subsystem that Yi has developed. Dave and Brian are looking into the PXI system for the solenoid.
- Dave will be working on the gas system controls. Namely he will implement the communications between PLC and the Brooks controllers to be able to set the most parameters in the controller downstairs from PLC. We would like to be able to change the flow or to completely shut off the valves in the MFC in the hall from the hall.
- Dave also suggested to have a NC solenoid valve for each FDC package so that we do not need to rely on the serial communications in case of emergency. It was pointed out that this is something that may trigger a review, the FDC group needs to make decisions on that.
Remote reset of CAENs
- Dave have plans for redoing the remote reset for CAENs without using 5V output modules that currently kill the HV from CAEN each time PLC is rebooted.
- This may require additional purchases of PLC equipment, Hovanes will ask Eugene about POA.
Warm magnet power supplies
- The procedures for TAG5C and SWEEP magnet ramping require approching the desired setpoint from above after staying near the saturation point for a while.
- During the fall run we had incidents for both PS and tagger where the magnet currents were not raised to the maximum values before putting them at the nominal setpoint.
- Accelerator magnets approach their desired setpoints from below after looping multiple times for degaussing, so the existing software does not accommodate TAG5C and SWEEP, but probably can be modified to execute our simple loop.
- Hovanes will talk to Gary about making such changes.
Solenoid controls work
- Mark would like to have the correct time maintained on the solenoid power supply. He thinks the precision might be limited to 1sec by the power supply, but he will be check that.
- Dave is planning to implement some mechanism by which the time will be transferred from PLC to the power supply either periodically or upon some condition being satisfied. He would like to have a PLC tag connected to EPICS that reports a correct time.
- Hovanes suggested to use the time from EPICS/PLC interface IOC for the solenoid. He will send Dave the name of the tag, and Dave will take care of the conversion. Hovanes will also see if the server on which the production version of that IOC is running can be synced with JLab time servers.
- Flow meters for the leads will be replaced by Scot.
- Dave is considering a possibility of being able to stop a slow dump and ramping PS back to the setpoint. This might not be so easy, if not impossible, to do without going to the fast dump, according to Mark.
- New liquid level sensor will be installed and needs to be integrated with the controls system.
- Monitoring circuit for GFD is being worked on by Mark.
- Mark will send Hovanes a list of LCW quantities which will be useful to have in the magnet section of the Hall D alarm system.
- Scot suggested to have the hall humidities and temperatures in the alarm system as well. Hovanes will add them.
- Two more motors will probably be installed for April running: beam profiler in front of the collimator and the polarimeter target in front of the east wall of the collimator cave.
- The specs for the polarimeter target motion system is still not well known, the details are coming in.
- The stage for the beam profiler is being designed, the cabling is expected to be very similar to collimator and harps.
- Target controls by the JLAB target group is pretty much done. They implemented the target controls in compact-RIO, but have an EPICS CA-server (or IOC) to interface it with Hall D EPICS. The target group needs a Windows PC in the counting house. we can share the halldsc7 that Dave is using for PLC with them. Hovanes will setup a meeting with the target group to work out the details of integration.
- Scot reminded us that TAC motor needs to be tested. It has been locked-out/tagged-out during the last fall running. Hovanes will get together and check that the TAC stage is fine.